# Bloom: Designing for LLM-Augmented Behavior Change Interactions

### v1

@@ 8. Limitations
Our study has several limitations. First, our study design was not intended to robustly evaluate behavior change in a clinical sense, which requires large-scale, randomized controlled trials conducted over several months to years. Consequently, our statistical tests were likely underpowered, and our study duration was likely insufficient for larger differences in PA habits to materialize. Since participants self-selected into the study and were compensated, the effect sizes observed here are likely larger than what would be expected in a naturalistic deployment. We concur with Klasnja et al. (2011), who have argued that evaluating behavior change in a traditional sense is often inappropriate for early-stage HCI systems research. Instead, our study’s main focus was on participant experiences through qualitative interviews and extensive surveys, aiming to uncover insights into why and how such systems succeed or fail. The four-week duration balanced high participation demands with known issues with sustained engagement in behavior change research (Yang et al., 2020). Further, our sample size is larger than is typical in previous HCI deployment studies targeting PA promotion (Consolvo et al., 2008; Murnane et al., 2023) or LLM health coaching (Xu et al., 2025).
Since our control condition did not include a rule-based chatbot, some of our findings may have also arisen with a non-LLM chatbot. However, our findings around relational cues and mindsets differ from prior work on rule-based coaches (Mitchell et al., 2021; Luo et al., 2021). Moreover, given that we did not observe significant differences in quantitative PA outcomes with an LLM chatbot, it is unlikely that a simpler chatbot would have produced stronger effects.
@@
Lastly, the Bloom system itself had technological limitations. We did not explore more advanced agent designs involving finetuning (Mantena et al., 2025; Khasentino et al., 2025), multi-agent systems (Merrill et al., 2024), or reasoning models (Jaech et al., 2024; Guo et al., 2025). While our summary-based memory module was well within the model’s context limit, even for users that chatted daily (¡3,500 tokens), more sophisticated memory systems (e.g., (Park et al., 2023)) could additionally have improved performance. Given the promising initial feedback we received on our prompt-based approach, we considered it sufficient for generating design insights. Moreover, advanced improvements would have required significant development and/or data collection efforts without clear evidence of user need or interaction requirements. We also did not explore all possible LLM-augmented behavior change interactions, such as social features or gamification. Nevertheless, we believe that qualitative context can be informative for a broad range of interactions and is a valuable exploration for future work.
@@ 9. Conclusion
In this work, we introduce Bloom, a mobile application integrating an LLM-based coaching chatbot with established behavior change interactions. Bloom leverages qualitative context from coaching conversations to further personalize LLM-augmented behavior change interactions, moving beyond existing text-only approaches in prior work on LLMs for behavioral health. Through a four-week field study involving 54 participants, we found that Bloom fostered positive mindsets toward PA and supported more personalized and flexible workout planning compared to a non-LLM control. Treatment participants reported greater improvements in psychological and motivational outcomes, including increased enjoyment, self-confidence, and a greater sense of agency in managing their PA goals. Both conditions substantially increased participants’ exercise levels, doubling the number of individuals who met or exceeded the recommended 150 min/week of PA. Although we observed no significant differences between conditions in objectively measured PA, psychological and motivational outcomes indicate the potential for sustained behavior change over longer periods when using our LLM-based health coach. Bloom presents a novel opportunity to create more supportive and empowering behavior change applications that leverage qualitative context to more sustainably shape people’s mindsets and motivation.
10. Acknowledgments
We are grateful for the funding support provided by the Stanford Institute for Human-Centered Artificial Intelligence (HAI) and the Hasso Plattner Foundation, as well as the OpenAI researcher access program for providing API credits to partially support this research. We thank the many members of the IxD and AI4HI research groups for their continuous support and feedback on this project, and all of our friends & family testers for catching bugs and suggesting improvements. We thank Michelle Lam, Lindsay Popowski, and Dulce Garcia for their assistance with participant recruitment; Ryan Louie, Ifdita Hasan, and Ankita Koodavoor for their advice on our redteaming and safety evaluation; Andrew Sung for his contributions to app development; and Carolyn Zhou for their feedback on the LLM agent architecture. We are also grateful to Tobias Gerstenberg, Nilam Ram, and Michael Bernstein for guidance on statistical modeling, to Elizabeth Murnane for her input on our ambient display design and validation study, and to Mary Czerwinski, Andrea Green, Omar Shaikh, and Helena Vasconcelos for thoughtful feedback on early drafts of this manuscript. Most importantly, we thank all of our participants, without whom this research would not have been possible.
@@ Appendix A LLM Coaching Agent
In this section, we provide further technical and implementation details on our LLM agent’s tools.
@@ A.1. Querying Health Data
The query_health_data function allows the agent to fetch the user’s wearable data leveraging Apple’s HealthKit API. This function is always available to the agent and has the following function signature:
@@
The HealthKit query is executed on-device in Swift and returns the aggregated data as text along with a natural language description of the data source to the LLM. If show_user=true, the data is displayed as a visualization in the chat (see Figure 3D). We encountered issues with LLMs being able to reliably perform date arithmetic, so we allow reference_date to be parsed either as a natural language string (e.g., “today”, “last week”) or a date string. Unlike GPTCoach (Jörke et al., 2025), we do not include a tool use chain (a prompt chain that forces the model to decide whether to augment its response with the user’s health data), which led to many unnecessary data queries and greatly increased latency.
@@ A.2. Structured Plan Generation
During onboarding and check-in chats, we expose the generate_plan function to the agent. This function generates a weekly plan using an external prompt that takes the current conversation history and the user’s plan history as input and produces a structured JSON object. After generating the plan as a JSON object, it is displayed to the user as a plan widget in the chat (see Figure 2D; Figure 3D) and saved to our database such that it can be rendered and modified in the app’s UI.
The plan generation prompt includes guidelines for creating a well-rounded, stage-appropriate, and personalized exercise plans sourced from the Active Choice program and the Center for Disease Control’s Physical Activity Guidelines (for Disease Control and Prevention, 2022). The agent is instructed to call the generate_plan function during the goal setting dialogue state. The preceding dialogue state prompts include explicit instructions to solicit relevant information and preferences from the user (e.g., goals, resources, injuries, past experience, preferred activities, and schedule), ensuring that the goal-setting process is collaborative, not prescriptive, and that the conversation history contains sufficient context for the generate_plan function to generate a personalized plan. We do not allow the agent to advance out of the goal setting dialogue state until the generate_plan function has been called. In absence of this manual check, we encountered hallucination issues with the agent telling the user it would generate a plan without calling the underlying function.
Moreover, we expose add_workout and delete_workout functions to the at-will chat agent, allowing the at-will agent to make direct edits to the user’s current plan via conversational interaction. We do not provide the at-will agent access to the generate_plan function—early tests indicated that hallucination and tool use errors could lead to highly destructive edits or deletions. Making small, granular edits to the current plan proved to be more reliable than generating a replacement plan directly.
@@ Appendix B Ambient Display Logic
Our ambient display encodings and logic were strongly inspired by UbiFit (Consolvo et al., 2008; Consolvo, 2008) and modified to match our four-week study progression. The ambient display advances (i.e., a flower in the garden grows) as the user completes workouts in their weekly plan in 20% increments. For example, a flower grows if completing a workout advances the plan progress from 30% to 40% completion, but not from 20% to 30%. The flower fully blooms when the weekly plan is 100% completed. If the user completes their plan in week $t$, the flower remains in the garden and a new flower starts growing in week $t+1$. If they do not complete their plan in week $t$, the flower starts growing again from scratch in week $t+1$. In addition to a fully-bloomed flower, we add additional elements (week 2: bird on a branch; week 3: beehive on a branch; week 4: bird and birdhouse) to the display.
For each completed workout, a critter is drawn to the ambient display above the growing flowers. Users receive a bee for walking and a butterfly for other activites, with different colored buttlerflies for different activity types (red: cardio, orange: strength, green: team sports, yellow: flexibility & dance, blue: outdoor recreation, purple: misc). A small (¡ 15min), medium (15-30 min), or large (30+ min) critter is drawn depending on the exercise duration. To avoid cluttering the display, critters are reset each week. After the final screen (week 4, 100%), the ambient display remains fixed in subsequent weeks, but critters continue to appear.
@@ Appendix C Field Study
In this section, we provide full details on our statistical analysis, custom survey questions, and complete survey results.
@@ C.1. Statistical Analysis
To account for systematic data missingness (e.g., due to privacy permissions, data upload errors, or participants forgetting to wear their watch), we excluded participants with fewer than 30 baseline days or with three or more missing study days from our analysis. This yielding a sample of 54 participants for both step count and distance walking/running (6,348 observations), 41 participants for energy burned (4,615 observations), and 36 participants for exercise time analyses (3,811 observations). Notably, relaxing our exclusion criteria and including all participants in our model does not change the directionality or significance of our effects.
@@
We explored a variety of models varying in complexity, from two-sample t-tests and standard repeated-measures ANOVA, to more sophisticated mixed-effects models including multiple levels, random slopes, and non-linear time trends. We also examined alternative coding schemes (e.g., treating weeks or days as factors) and aggregating data at the weekly rather than daily level. We selected this final model because it provided the best statistical fit to the data (lowest AIC/BIC), leveraged daily-level observations to maximize statistical power, and avoided the strong assumption that treatment effects vary linearly at the daily level. Instead, the model captures treatment effects as linear weekly trends, aligning well with our hypothesis concerning persistence (H3). Additionally, this approach explicitly accounts for heteroskedasticity by allowing participant-specific residual variances, acknowledging that observed changes in activity are more noteworthy for participants whose baseline behavior was less variable. Importantly, the patterns of significance, directionality of effects, and substantive conclusions drawn from our results remained consistent across all modeling approaches tested.
C.2. Survey Measures
@@
    During the Physical Activity st 7 days, how much exercise did you get?
@@
Stage of Change (Marcus et al., 1992) (1: pre-contemplation, 5: maintenance) 1.88 (0.59) 2.85 (0.73) 0.96 (0.96) 1.89 (0.74) 2.61 (1.07) 0.71 (1.01) 0.25
IPAQ (Craig et al., 2003) (MET‑min) 415.31 (329.00) 749.73 (467.88) 334.42 (505.20) 548.18 (637.10) 893.96 (741.87) 345.79 (965.50) ‑11.36
@@
Self‑Efficacy (Sallis et al., 1988) 2.83 (0.67) 2.82 (0.88) ‑0.01 (0.93) 2.61 (0.76) 2.63 (0.86) 0.02 (0.99) ‑0.03
@@
Barriers to Being Active (Rodríguez et al., 2008) 2.59 (0.56) 2.35 (0.63) ‑0.24 (0.36) 2.64 (0.62) 2.38 (0.71) ‑0.27 (0.64) 0.02
SUS (Brooke et al., 1996) (0–100) 93.85 (8.89) 82.21 (20.85) ‑11.63 (18.64) 93.75 (8.29) 89.02 (8.91) ‑4.73 (8.80) ‑6.90
SASSI (Hone and Graham, 2000) 4.40 (0.43) 4.07 (0.75) ‑0.33 (0.63) 4.38 (0.36) 4.25 (0.44) ‑0.13 (0.46) ‑0.20
TSRI (Bentvelzen et al., 2021) 3.80 (0.58) 3.85 (0.86) 0.06 (0.77) 3.93 (0.53) 3.61 (0.92) ‑0.33 (0.80) 0.38
eHealth Literacy (Norman and Skinner, 2006) 4.46 (0.50) 4.35 (0.82) ‑0.11 (0.80) 4.44 (0.53) 4.59 (0.52) 0.15 (0.39) ‑0.26
(Adapted) SDM-Q-9 (Kriston et al., 2010) 4.06 (0.64) 4.18 (0.79) 0.12 (0.46) 3.75 (0.85) 3.48 (0.92) ‑0.26 (0.69) 0.38
@@
Question Mean Day 1 Day 28 $\Delta_{\text{T}}$ $\beta_{\text{T}}$ Mean Day 1 Day 28 $\Delta_{\text{C}}$ $\beta_{\text{C}}$
I am satisfied with my current level of physical activity. 3.17 (1.09) 2.59 (1.18) 3.45 (1.01) 1.17 (0.99) 0.022 3.06 (1.07) 2.71 (1.10) 2.91 (1.16) 0.43 (1.22) 0.006
I am a physically active person. 3.15 (1.03) 2.77 (1.02) 3.32 (0.95) 0.61 (0.78) 0.030 3.02 (1.05) 2.53 (0.87) 3.00 (1.00) 0.71 (0.99) 0.009
I am committed to my physical activity goal. 4.27 (0.66) 4.23 (0.53) 4.09 (0.81) 0.17 (0.38) -0.007 3.96 (0.84) 3.82 (0.73) 3.91 (0.90) -0.21 (1.31) 0.000
How hopeful are you about your physical health today? 4.07 (0.70) 3.91 (0.68) 4.14 (0.71) 0.44 (0.78) 0.005 3.84 (0.85) 3.88 (0.60) 3.87 (0.87) -0.21 (0.80) -0.003
How was your mood today? (unpleasant/pleasant) 3.93 (0.80) 4.14 (0.64) 4.32 (0.84) 0.33 (0.91) 0.004 3.91 (0.83) 4.00 (0.71) 3.96 (0.77) -0.21 (0.97) 0.002
Table 10. Daily Survey Results. We report mean (SD) values by condition across all days, for day 1, and for day 28. $\Delta$ is the change from day 1 to day 28 (mean difference, SD). $\beta$ is the per-day slope from a linear regression of score on day. All responses map to a 5 point Likert scale.
@@
Question Mean Week 1 Week 4 $\Delta_{\text{T}}$ $\beta_{\text{T}}$ Mean Week 1 Week 4 $\Delta_{\text{C}}$ $\beta_{\text{C}}$
Overall App Rating 3.84 (0.74) 3.90 (0.67) 3.98 (0.81) -0.05 (0.63) -0.01 4.02 (0.69) 3.94 (0.68) 4.02 (0.68) 0.30 (0.69) 0.04
Visual Wallpaper Rating 3.78 (1.12) 3.70 (1.18) 3.93 (1.22) 0.29 (0.61) 0.08 3.94 (0.94) 3.73 (0.98) 4.14 (0.81) 0.24 (0.82) 0.15
Exercising is… (difficult/easy) 3.11 (0.82) 2.92 (0.78) 3.35 (1.11) 0.39 (1.03) 0.14 2.93 (1.00) 2.90 (1.17) 2.87 (0.98) 0.00 (0.50) 0.02
Exercising is… (boring/fun) 3.17 (0.91) 3.13 (0.86) 3.27 (1.03) 0.24 (0.70) 0.04 2.96 (1.06) 2.81 (1.10) 2.87 (1.10) 0.09 (0.94) 0.06
Barriers To Being Active (4 item subset) 2.54 (0.70) 2.51 (0.74) 2.53 (0.71) -0.02 (0.66) -0.01 2.79 (0.85) 2.92 (0.86) 2.60 (0.96) -0.44 (0.39) -0.06
Table 11. Weekly Survey Results. We report mean (SD) values by condition across all weeks, for week 1, and for week 4. $\Delta$ is the change from week 1 to week 4 (mean difference, SD). $\beta$ is the per-week slope from a linear regression of score on week. All responses were rescaled to map to a 5 point Likert scale.

### v2

@@ 8. Limitations
Our study has several limitations. First, our study design was not intended to robustly evaluate behavior change in a clinical sense, which requires large-scale, randomized controlled trials conducted over several months to years. Consequently, H2 and H3 were underpowered to detect moderate effect sizes, and our study duration was likely insufficient for larger differences in PA habits to materialize. Since participants self-selected into the study and were compensated, the effect sizes observed here are likely larger than what would be expected in a naturalistic deployment. As argued by Klasnja et al. (2011), evaluating behavior change in a traditional sense is often inappropriate for early-stage HCI systems research. Instead, our study’s main focus was on studying participant experiences through qualitative interviews and extensive surveys, aiming to uncover insights into why and how such systems succeed or fail. The four-week duration balanced high participation demands with known issues with sustained engagement in behavior change research (Yang et al., 2020).
Since our control condition did not include a rule-based chatbot, some of our findings may have also arisen with a non-LLM chatbot. However, our findings around relational cues and mindsets differ from prior work on rule-based coaches (Mitchell et al., 2021; Luo et al., 2021). Moreover, we consider it unlikely that a simpler chatbot would have produced stronger effects on quantitative PA outcomes.
@@
Lastly, the Bloom system itself had technological limitations. We did not explore more advanced agent designs involving finetuning (Mantena et al., 2025; Khasentino et al., 2025), multi-agent systems (Merrill et al., 2026), or reasoning models (Jaech et al., 2024; Guo et al., 2025). While our summary-based memory module was well within the model’s context limit, even for users that chatted daily (¡3,500 tokens), more sophisticated memory systems (e.g., (Park et al., 2023)) could have improved performance. Given the promising initial feedback we received on our prompt-based approach, we considered it sufficient for generating design insights. Moreover, advanced improvements would have required significant development and/or data collection efforts without clear evidence of user need or interaction requirements. We also did not explore all possible LLM-augmented behavior change interactions, such as social features or gamification.
@@ 9. Conclusion
In this work, we introduce Bloom, a mobile application integrating an LLM-based coaching chatbot with established behavior change interactions. Bloom leverages qualitative context from coaching conversations to further personalize LLM-augmented behavior change interactions, moving beyond existing text-only approaches in prior work on LLMs for behavioral health. Through a four-week field study involving 54 participants, we found that Bloom fostered positive mindsets toward PA and supported more personalized and flexible workout planning compared to a no-LLM control. Treatment participants reported greater improvements in mindset-related outcomes, including increased enjoyment, self-confidence, and a greater sense of agency in managing their PA goals. Both conditions substantially increased participants’ exercise levels, doubling the number of individuals who met or exceeded the recommended 150 min/week of PA. Although we observed no advantage for the LLM condition in short-term PA levels, psychological and motivational outcomes indicate the potential for sustained behavior change over longer periods when using our LLM-based health coach. Bloom presents a novel opportunity to create more supportive and empowering behavior change applications that leverage qualitative context to more sustainably shape people’s mindsets and motivation.
Acknowledgements.
We are grateful for the funding support provided by the Stanford Institute for Human-Centered Artificial Intelligence (HAI), the Hasso Plattner Foundation, Hanwha, Wellhub, as well as the OpenAI researcher access program for providing API credits to partially support this research. We thank the many members of the IxD and AI4HI research groups for their continuous support and feedback on this project, and all of our friends & family testers for catching bugs and suggesting improvements. We thank Michelle Lam, Lindsay Popowski, and Dulce Garcia for their assistance with participant recruitment; Ryan Louie, Ifdita Hasan, and Ankita Koodavoor for their advice on our redteaming and safety evaluation; Andrew Sung for his contributions to app development; and Carolyn Zhou for their feedback on the LLM agent architecture. We are also grateful to Tobias Gerstenberg, Nilam Ram, and Michael Bernstein for guidance on statistical modeling; to Elizabeth Murnane for her input on our ambient display design and validation study; to Pedja Klasnja for his advice on post hoc power analyses; and to Mary Czerwinski, Andrea Green, Omar Shaikh, and Helena Vasconcelos for thoughtful feedback on early drafts of this manuscript. Most importantly, we thank all of our participants, without whom this research would not have been possible.
Supplementary Materials
@@ Appendix A LLM Coaching Agent
In this section, we provide further technical and implementation details about our LLM agent’s tools.
@@ A.1. Querying Health Data
The query_health_data function allows the agent to fetch the user’s wearable data using Apple’s HealthKit API. This function is always available to the agent and has the following function signature:
@@
The HealthKit query is executed on device in Swift and returns the aggregated data as text along with a natural language description of the data source to the LLM. If show_user=true, the data is displayed as a visualization in the chat (see Figure 3D). We encountered issues with LLMs’ ability to reliably perform date arithmetic, so we allow reference_date to be parsed either as a natural language string (e.g., “today”, “last week”) or a date string. Unlike GPTCoach (Jörke et al., 2025), we do not include a tool use chain (a prompt chain that forces the model to decide whether to augment its response with the user’s health data), which led to many unnecessary data queries and greatly increased latency.
@@ A.2. Structured Plan Generation
During onboarding and check-in chats, we expose the generate_plan function to the agent. This function generates a weekly plan using an external prompt that takes the current conversation history and the user’s plan history as input and produces a structured JSON object. After generating the plan as JSON, it is displayed to the user as a plan widget in the chat (see Figure 2D; Figure 3D) and saved to our database such that it can be rendered and modified in the app’s UI. The plan generation JSON schema and prompt are provided in our GitHub repository.
The plan generation prompt includes guidelines for creating a well-rounded, stage-appropriate, and personalized exercise plan sourced from the Active Choices program and the Centers for Disease Control and Prevention’s Physical Activity Guidelines (for Disease Control and Prevention, 2022). The agent is instructed to call the generate_plan function during the goal setting dialogue state. The preceding dialogue state prompts include explicit instructions to solicit relevant information and preferences from the user (e.g., goals, resources, injuries, past experience, preferred activities, schedule), ensuring that the goal setting process is collaborative, not prescriptive, and that the conversation history contains sufficient context for the generate_plan function to generate a personalized plan. We do not allow the agent to advance out of the goal setting dialogue state until the generate_plan function has been called. In the absence of this manual check, we encountered hallucination issues with the agent telling the user it would generate a plan without calling the underlying function.
Moreover, we expose add_workout and delete_workout functions to the at-will chat agent, allowing the at-will agent to make direct edits to the user’s current plan via conversational interaction. We do not provide the at-will agent with access to the generate_plan function—early tests indicated that hallucination and tool use errors could lead to highly destructive edits or deletions. Making small, granular edits to the current plan proved to be more reliable than generating a replacement plan directly.
@@ Appendix B Ambient Display Logic
Our ambient display encodings and logic were strongly inspired by UbiFit (Consolvo et al., 2008; Consolvo, 2008) and modified to match our four-week study progression. The ambient display advances (i.e., a flower in the garden grows) as the user completes workouts in their weekly plan in 20% increments. For example, a flower grows if completing a workout advances the plan progress from 30% to 40% completion, but not from 20% to 30%. The flower fully blooms when the weekly plan is 100% completed. If the user completes their plan in week $t$, the flower remains in the garden and a new flower starts growing in week $t+1$. If they do not complete their plan in week $t$, the flower starts growing again from scratch in week $t+1$. In addition to a fully bloomed flower, we add additional elements (week 2: bird on a branch; week 3: beehive on a branch; week 4: bird and birdhouse) to the display.
For each completed workout, a critter is drawn to the ambient display above the growing flowers. Users receive a bee for walking and a butterfly for other activities, with different colored butterflies for different activity types (red: cardio, orange: strength, green: team sports, yellow: flexibility & dance, blue: outdoor recreation, purple: misc). A small (¡ 15min), medium (15-30 min), or large (30+ min) critter is drawn depending on the exercise duration. To avoid cluttering the display, critters are reset each week. After the final screen (week 4, 100%), the ambient display remains fixed in subsequent weeks, but critters continue to appear.
@@ C.1. Statistical Analysis
To account for systematic data missingness (e.g., due to privacy permissions, data upload errors, or participants forgetting to wear their watch), we excluded participants with fewer than 30 baseline days or with three or more missing study days from our analysis. This yielded a sample of 54 participants for both step count and distance walking/running (6,348 observations), 41 participants for energy burned (4,615 observations), and 36 participants for exercise time analyses (3,811 observations). Notably, relaxing our exclusion criteria and including all participants in our model does not change the directionality or significance of our effects.
@@
We explored a variety of models varying in complexity, from two-sample t-tests and standard repeated-measures ANOVA, to more sophisticated mixed-effects models including multiple levels, random slopes, and nonlinear time trends. We also examined alternative coding schemes (e.g., treating weeks or days as factors) and aggregating data at the weekly rather than daily level. We selected this final model because it provided the best statistical fit to the data (lowest AIC and BIC), leveraged daily observations to maximize statistical power, and avoided the strong assumption that treatment effects vary linearly at the daily level. Instead, the model captures treatment effects as linear weekly trends, aligning well with our hypothesis concerning persistence (H3). Additionally, this approach explicitly accounts for heteroskedasticity by allowing participant-specific residual variances, acknowledging that observed changes in activity are more noteworthy for participants whose baseline behavior was less variable. Importantly, the patterns of significance, directionality of effects, and substantive conclusions drawn from our results remained consistent across all modeling approaches tested.
C.2. Power Considerations
Given the exploratory nature of our study, we did not conduct a formal a priori power analysis. To aid in the interpretation of our quantitative results, we performed several post hoc analyses. First, we conduct a sensitivity analysis to estimate the sample sizes required to achieve 80% power at a significance level of $\alpha=.05$ for a range of plausible effect sizes. We consider a small effect ($d=0.3$) commonly observed in meta-analyses of mHealth interventions (Yang and Van Stee, 2019; Laranjo et al., 2021), as well as medium ($d=0.5)$ and large ($d=0.8$) effects that might be expected in early-stage work. Second, given our actual sample size ($N=54$, assuming an equal split across conditions), we compute the minimally detectable effect size.
We use standard $t$-test approximations for both analyses because a full simulation of a three-level mixed-effects model would require specifying a complete generative model for all variance components, covariance structures, and missingness patterns. Instead, we opted for more transparent calculations that rely on fewer assumptions, treating H1 as a paired $t$-test and H2 as a two-sample $t$-test. This approximation provides an upper bound on the required sample sizes, as our actual mixed-effects model leverages repeated daily observations and partial pooling across weeks, which typically increases statistical efficiency. We exclude H3 because power for slope-difference interactions would similarly require full simulations of the multilevel model, and, given that higher-order interactions are generally harder to detect than mean differences (H2), would be expected to require an even larger sample. We use the R pwr package for our calculations.
Results indicate that detecting a small effect ($d=0.3$) would require at least 90 participants for H1 and at least 352 participants for H2. A moderate effect ($d=0.5$) would require at least 34 participants for H1 and at least 128 participants for H2. A large effect ($d=0.8$) would require at least 15 participants for H1 and at least 52 participants for H2. Given our actual sample ($N=54$), the minimally detectable effect sizes were $d\geq 0.39$ for H1 and $d\geq 0.78$ for H2. Thus, the study had adequate power to detect moderate or large within-person improvements (H1) but was underpowered to detect small or moderate between-condition effects (H2/H3). We observed consistently large within-person improvements for H1 (d=0.59d=0.590.750.75) and small between-condition and persistence effects for H2 and H3 (including small negative effects, d=0.22d=-0.220.230.23), indicating that H1 had sufficient power but H2 and H3 were underpowered.
C.3. Supplemental Analysis: Decomposing H1 into Within-Condition Effects
In the main text, H1 is evaluated as an aggregate effect across both conditions, testing whether the study period showed meaningful improvements over the pre-study baseline regardless of condition. In this section, we decompose H1 into separate within-condition tests (H1a/H1b) to assess whether each group shows meaningful pre-post improvement.
  • H1a:
    Mean PA levels in the control group will exceed their baseline levels during the study period.
  • H1b:
    Mean PA levels in the treatment group will exceed their baseline levels during the study period.
  • H2:
    The treatment group’s increase in mean PA levels from baseline to the study period will exceed that of the control group.
  • H3:
    The treatment group will exhibit a smaller rate of decline (or greater persistence) in PA levels across the four-week study period compared to the control group.
This decomposition allows H1b to provide early-stage evidence of a meaningful improvement in the treatment group, analogous to proof-of-concept evidence in the ORBIT model (Czajkowski et al., 2015). We emphasize, however, that our study was not designed as a proof-of-concept trial and its primary aim was to surface design insights. We did not revise the hypothesis structure of the main text, as this framing reflects a post hoc interpretive consideration rather than the basis for our original study design.
These tests were conducted using the same three-level linear mixed-effects model described in C.1, except that H1a and H1b were fit separately within each condition with treatment interaction terms ($\beta_{2}$ and $\beta_{4}$) removed. In Table 9, we report coefficient means, standard errors, and significance levels for each hypothesis, analogous to Table 7 in the main text. Across all four PA outcomes (step count, energy burned, distance walked/run, and exercise time), both H1a and H1b were significant ($p<0.01$), with meaningful improvements in the treatment condition (e.g., +994 steps/day or 81.9 minutes of exercise per week).
PA Outcome Hyp. Description Estimate (SE)
Step Count H1a Difference in daily step count from baseline to study period (control group) +1 648 (281)***
H1b Difference in daily step count from baseline to study period (treatment group) +994 (337)**
Active Energy H1a Difference in daily kcal burned from baseline to study period (control group) +87.8 (23.9)***
Burned (kcal) H1b Difference in daily kcal burned from baseline to study period (treatment group) +79.9 (20.9)***
H2 Treatment–control difference in daily kcal burned change from baseline to study period –7.09 (31.3)
Exercise Time (min) H1a Difference in daily exercise min from baseline to study period (control group) +12.6 (3.74)**
H1b Difference in daily exercise min from baseline to study period (treatment group) +11.7 (3.49)**
Distance Walking/ H1a Difference in daily distance walked/run (km) from baseline to study period (control group) +0.753 (0.139)***
Running (km) H1b Difference in daily distance walked/run (km) from baseline to study period (treatment group) +0.519 (0.144)***
H2 Treatment–control difference in daily distance walked/run (km) change from baseline to study period –0.239 (0.198)
H3 Treatment–control difference in the weekly rate of change ($\Delta$ km/day per study week) +0.0712 (0.0625)
Table 9. Within-Condition Wearable Data (Quantitative) Results. We report mean (SE) parameter estimates from models fit separately within each condition (H1a/b) and from the combined model (H2/H3). *** denotes $p<0.001$ and ** denotes $p<0.01$, where $p$-values are Holm-adjusted for multiple comparisons. Significant results are bolded.
C.4. Survey Measures
@@
    During the last 7 days, how much exercise did you get?
@@
Stage of Change (Marcus et al., 1992) (1: pre-con., 5: maint.) 1.88 (0.59) 2.85 (0.73) 0.96 (0.96) 1.89 (0.74) 2.61 (1.07) 0.71 (1.01) 0.25
IPAQ (Craig et al., 2003) (MET-min) 415.31 (329.00) 749.73 (467.88) 334.42 (505.20) 548.18 (637.10) 893.96 (741.87) 345.79 (965.50) -11.36
@@
Self-Efficacy (Sallis et al., 1988) 2.83 (0.67) 2.82 (0.88) -0.01 (0.93) 2.61 (0.76) 2.63 (0.86) 0.02 (0.99) -0.03
@@
Barriers to Being Active (Centers for Disease Control and Prevention, 2022) 2.59 (0.56) 2.35 (0.63) -0.24 (0.36) 2.64 (0.62) 2.38 (0.71) -0.27 (0.64) 0.02
SUS (Brooke and others, 1996) (0–100) 93.85 (8.89) 82.21 (20.85) -11.63 (18.64) 93.75 (8.29) 89.02 (8.91) -4.73 (8.80) -6.90
SASSI (Hone and Graham, 2000) 4.40 (0.43) 4.07 (0.75) -0.33 (0.63) 4.38 (0.36) 4.25 (0.44) -0.13 (0.46) -0.20
TSRI (Bentvelzen et al., 2021) 3.80 (0.58) 3.85 (0.86) 0.06 (0.77) 3.93 (0.53) 3.61 (0.92) -0.33 (0.80) 0.38
eHealth Literacy (Norman and Skinner, 2006) 4.46 (0.50) 4.35 (0.82) -0.11 (0.80) 4.44 (0.53) 4.59 (0.52) 0.15 (0.39) -0.26
(Adapted) SDM-Q-9 (Kriston et al., 2010) 4.06 (0.64) 4.18 (0.79) 0.12 (0.46) 3.75 (0.85) 3.48 (0.92) -0.26 (0.69) 0.38
@@
Question Mean Day 1 Day 28 $\beta_{\text{T}}$ $\Delta_{\text{T}}$ Mean Day 1 Day 28 $\beta_{\text{C}}$ $\Delta_{\text{C}}$
I am satisfied with my current level of physical activity. 3.17 (1.09) 2.59 (1.18) 3.45 (1.01) 0.022 0.62 3.06 (1.07) 2.71 (1.10) 2.91 (1.16) 0.006 0.17
I am a physically active person. 3.15 (1.03) 2.77 (1.02) 3.32 (0.95) 0.030 0.84 3.02 (1.05) 2.53 (0.87) 3.00 (1.00) 0.009 0.25
I am committed to my physical activity goal. 4.27 (0.66) 4.23 (0.53) 4.09 (0.81) -0.007 -0.21 3.96 (0.84) 3.82 (0.73) 3.91 (0.90) -0.000 -0.01
How hopeful are you about your physical health today? 4.07 (0.70) 3.91 (0.68) 4.14 (0.71) 0.005 0.14 3.84 (0.85) 3.88 (0.60) 3.87 (0.87) -0.003 -0.09
How was your mood today? (unpleasant/pleasant) 3.93 (0.80) 4.14 (0.64) 4.32 (0.84) 0.004 0.12 3.91 (0.83) 4.00 (0.71) 3.96 (0.77) 0.002 0.06
Table 12. Daily Survey Results. We report mean (SD) values by condition across all days, for day 1, and for day 28. $\beta$ is the per-day slope from a linear regression of score on day. $\Delta$ is the predicted 28-day change computed as $28\times\beta$. All responses map to a 5-point Likert scale.
@@
Question Mean Week 1 Week 4 $\beta_{\text{T}}$ $\Delta_{\text{T}}$ Mean Week 1 Week 4 $\beta_{\text{C}}$ $\Delta_{\text{C}}$
Overall App Rating 3.84 (0.74) 3.90 (0.67) 3.98 (0.81) -0.005 -0.02 4.02 (0.69) 3.94 (0.68) 4.02 (0.68) 0.040 0.16
Visual Wallpaper Rating 3.78 (1.12) 3.70 (1.18) 3.93 (1.22) 0.077 0.31 3.94 (0.94) 3.73 (0.98) 4.14 (0.81) 0.153 0.61
Exercising is … (difficult/easy) 3.11 (0.82) 2.92 (0.78) 3.35 (1.11) 0.139 0.56 2.93 (1.00) 2.90 (1.17) 2.87 (0.98) 0.024 0.10
Exercising is … (boring/fun) 3.17 (0.91) 3.13 (0.86) 3.27 (1.03) 0.045 0.18 2.96 (1.06) 2.81 (1.10) 2.87 (1.10) 0.056 0.22
Barriers To Being Active (4 item subset) 2.54 (0.70) 2.51 (0.74) 2.53 (0.71) -0.009 -0.04 2.79 (0.85) 2.92 (0.86) 2.60 (0.96) -0.059 -0.24
Table 13. Weekly Survey Results. We report mean (SD) values by condition across all weeks, for week 1, and for week 4. $\beta$ is the per-week slope from a linear regression of score on week. $\Delta$ is the predicted four-week change, computed as $4\times\beta$. All responses were rescaled to a 5-point Likert scale.